OpenStudy (anonymous):

f(x) = 2x^2 − 4x − 15 what are the x-intercepts and vertex point?

OpenStudy (mimi_x3):

x-intercept: y=0 2x^2 - 4x - 15 = 0 x = 3.91, -1.91

OpenStudy (anonymous):

i'm stuck on this problem on an online homework assignment. it didn't accept that answer.

OpenStudy (mimi_x3):

Would it accept the answer in surd form ?

OpenStudy (anonymous):

i don't know unless i try lol

OpenStudy (mimi_x3):

\[1/2(2-\sqrt{34})\] or \[1/2(2+\sqrt{34})\]

OpenStudy (anonymous):

it worked, thank you so much. so many people looked at my question and left lol, thank so for answering

OpenStudy (mimi_x3):

yay ! it worked (: yw (: